Question: A die is rolled and a spiner (4 sides) is spun. Find the probability:
P ( an even number and C)
P ( 2 or 5 and B or D)
P ( a number less than 5 and B or D)
One card is drawn from a standard deck of 52 cards. Find each probability
P( Jack or ten)
P(queen or Club)
( (diamond or black)

Answer by stanbon(75887)

Question: A die is rolled and a spiner (4 sides) is spun. Find the probability:
P ( an even number and C) = (3/6)*(1/4)
P ( 2 or 5 and B or D) = (2/6)*(2/4)
P ( a number less than 5 and B or D)= (4/6)*(2/4)
One card is drawn from a standard deck of 52 cards. Find each probability
P( Jack or ten)= (4/52) + (4/52)
P(queen or Club) = (4/52) + (13/52) - (1/52)
( (diamond or black)= (13/52) + (26/52)
==============================================
Cheers,
Stan H.